Calculate the molar mass of (CH3) 2CO. a. 58 g 18. b. 42 g c. 62 g d. 52 g e. 54 g Calculate the percentage by mass of oxygen in potassium permanganate, KMnO4. a. 24.7% 19. b. 34.7% с. 40.5% d. 32.290 38.6% e.
Answer #1

18) molar mass of (CH3)2CO:

=2 (molar mass of C) + 1 (molar mass of oxygen) + 6 (molar mass of hydrogen)

= 3(12.0107) + 1 (16.00) + 6 (1.00794)

= 36.0 + 16.0 + 6.05 = 58 g

Therefore, the molar mass of (CH3)2CO = 58 g

option A is correct

19) molar mass of KMnO4 = 158.034 g/mol

percentage by mass of oxygen = (mass of oxygen / moalr mass of KMnO4)*100

percentage by mass of oxygen = (64 / 158.034)*100 = 40.5 %

Therefore, option C is correct